Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

experts help plz

Expert replies
by quantskillsgmat » Sat Mar 10, 2012 9:55 pm
Q two identical urns-black and white each contain 5 blue,5 red and 10 green balls.every ball selected from black urn is immediately returned to the urn,while each ball selected from white urn is removed and placed on a table.If jenny receives a quarter for every blue ball,a dime for every red ball and a nickel for every green ball she selects.what is probability that she will be able to buy a 25 cent candy bar with the proceeds from drawing four balls two from each urn.
a)143/152 b)14/37 c)1 d)0 e)2/3
Join the discussion
Source: — Problem Solving |

by Snowbirds » Sun Mar 11, 2012 2:59 am
The only scenario where Jenny can't purchase the candy bar is if she selects 4 green balls. Therefore, overall probability is 1-P(4 green balls)

From the black urn:
1/2 * 1/2 = 1/4

From the white urn:
1/2 * 9/19 = 9/38

Combining together (1/4)*(9/38) = 9/152

1-9/152 = 143/152

a
Join the discussion

by factor26 » Sun Mar 11, 2012 7:14 am
Very nice post snowbird - it is also important to stress here that we should be multiplying the undesired outcomes (prob of 2 green from black * prob of 2 green from white) as the only way to not afford a candy bar in this situation is by drawing 2 green from one urn AND 2 green from the other. Additing them would indicate both situations are possible in obtaining our result.
Join the discussion